# Mapping class groups with the twisted Rokhlin property

### 05 Dec 2023 v1

@@ Abstract.
In this paper, generalising the idea of the Rokhlin property, we explore the concept of the twisted Rokhlin property of topological groups. A topological group is said to exhibit the twisted Rokhlin property if, for each automorphism $\phi$ of the group, there exists a $\phi$-twisted conjugacy class that is dense in the group. We provide a complete classification of connected orientable infinite-type surfaces without boundaries whose mapping class groups possess the twisted Rokhlin property. Additionally, we prove that the mapping class groups of the remaining surfaces do not admit any dense $\phi$-twisted conjugacy class for any automorphism $\phi$. This supplements the recent work of Lanier and Vlamis on the Rokhlin property of big mapping class groups.
@@
Big mapping class group; infinite type surface; Rokhlin property; twisted conjugacy; twisted Rokhlin property
@@ 1. Introduction
Given an automorphism $\phi$ of a group $G$, the map $(g,h)\mapsto gh\phi(g)^{-1}$, defines a left action of $G$ on itself. Two elements $x,y\in G$ are said to be $\phi$-twisted conjugate if they are in the same orbit under this action. The resulting orbits are referred as $\phi$-twisted conjugacy classes. The idea of twisted conjugacy arose from the work of Reidemeister [18]. It has deep connection with Nielsen fixed-point theory [5], and appears in Arthur–Selberg theory [20], algebraic geometry [9], and representation theory [7], to name a few. In recent years, there has been a tremendous amount of work on understanding twisted conjugacy in various classes of groups. See, for example [6], for a recent work on twisted conjugacy classes in mapping class groups and braid groups.
@@ 2.1. Classification of surfaces
An end is called planar if it has a neighbourhood that can be embedded in the plane. Otherwise, it is called non-planar. Let $\mathcal{E}_{np}(S)$ be the subspace of $\mathcal{E}(S)$ consisting of non-planar ends. Clearly, $\mathcal{E}_{np}(S)$ is a closed set in $\mathcal{E}(S)$. The following theorem gives a classification of surfaces that we shall need [19, Theorem 1].
@@
Without loss of generality, we shall assume throughout that the surface $S$ is symmetric about the $xy$-plane when embedded in $\mathbb{R}^{3}$. Since $\mathcal{E}(S)$ is homeomorphic to a closed subset of the Cantor set, it can be embedded on the great circle of intersection of the $xy$-plane and the 2-sphere $\mathbb{S}^{2}$. Now, take a sequence of disjoint closed disks $D_{j}$ on $\mathbb{S}^{2}\setminus\mathcal{E}(S)$ which are symmetric about the $xy$-plane such that the set of accumulation points of this sequence is precisely $\mathcal{E}_{np}(S)$. We remove the interiors of the disks $D_{j}$ and suitably identify in pairs the boundary circles by orientation-reversing homeomorphisms such that the desired resulting surface is again symmetric about the $xy$-plane. See [19, Theorem 3] for the general construction. For instance, Figure 1 illustrates the construction of Loch Ness monster surface, where the infinite sequence of disjoint closed disks $\{D_{i}\}_{i\in\mathbb{N}}$ is chosen such that the radii of the disks decreases as it approaches a point of $\mathcal{E}(S)$.
@@
For a connected orientable surface $S$ (not necessarility of infinite-type), let $\operatorname{Homeo}^{+}(S)$ denote the group of orientation-preserving self-homeomorphisms of $S$. We define a binary relation $\preceq$ on the space of ends $\mathcal{E}(S)$ of $S$. For $x,y\in\mathcal{E}(S)$, we say that $y\preceq x$ if for every open neighbourhood $\mathcal{U}$ of $x$, there is an open neighbourhood $\mathcal{V}$ of $y$ and a homeomorphism $h\in\operatorname{Homeo}^{+}(S)$ such that $h(\mathcal{V})\subseteq\mathcal{U}$. An end $\mu$ is called maximal if, for any end $e$ satisfying $\mu\preceq e$, we must have $e\preceq\mu$. Let $\mathcal{M}(S)$ denote the set of all maximal ends of $S$. Note that for every end $e\in\mathcal{E}(S)$, there exists a maximal end $\mu\in\mathcal{M}(S)$ such that $e\preceq\mu$ [15, Proposition 4.7]. We say that two ends $e_{1}$ and $e_{2}$ are comparable if either $e_{1}\preceq e_{2}$ or $e_{2}\preceq e_{1}$. The space $\mathcal{E}(S)$ is said to be self-similar if given a decomposition $\mathcal{E}(S)=\mathcal{E}_{1}\sqcup\cdots\sqcup\mathcal{E}_{n}$ into a finite disjoint union of clopen subsets, there exists an index $i\in\{1,\ldots,n\}$ such that $\mathcal{E}_{i}$ contains an open homeomorphic copy of $\mathcal{E}(S)$. The idea of maximality and self-similarity for end spaces is due to [15].
@@ 2.2. Mapping class groups of surfaces
For a simple closed curve $c$ on a surface $S$, let $T_{c}$ denote the left-hand Dehn twist along $c$. By abuse of notation, we denote a curve and its isotopy class by the same symbol, and use the same convention for homeomorphisms as well. The following lemma, perhaps well-known, gives a relation between the Dehn twist and its conjugate by a homeomorphism of $S$. The proof of the following lemma is a slight modification of the arguments of [4, Fact 3.7]. See also [12, Lemma 2.1].
@@
is a basis for the topology on $\mathrm{MCG}(S)$. Note that, the center of the big mapping class group is always trivial [13, Proposition 2]. The following basic observation shall be used later.
@@ 2.3. Curve graphs of surfaces
We say that a simple closed curve $\alpha$ on a surface $S$ is non-separating if $S\setminus\alpha$ is connected; otherwise, it is called separating. The intersection number $i(\alpha,\beta)$ between two curves $\alpha,\beta$ on $S$ is defined to be the minimum number of points of intersections between the representatives in their isotopy classes. A multicurve on $S$ is a set of pairwise (non-isotopic) disjoint curves.
The set $C(S)$ can be viewed as a graph whose vertex set is $C(S)$ and there is an edge between two vertices if the corresponding curves on $S$ are disjoint. The graph $C(S)$ is referred as curve graph of $S$. There is a natural metric $d_{C(S)}$ on $C(S)$ by defining the length of each edge as 1. We shall need the following result [16, Proposition 4.6].
@@
[17, Lemma 2.2] If $c\in\mathcal{C}(S)$ and $b,b^{\prime}\in\mathcal{C}(\Sigma)$ are two projections of $c$, then $d_{\mathcal{C}(\Sigma)}\left(b,b^{\prime}\right)\leq 2$.
@@ 3. Twisted Rokhlin property
    The homeomorphism group of the Cantor set [8, Theorem 2.6].
@@
    The group of orientation-preserving homeomorphisms of an even-dimensional sphere [8, Theorem 3.4].
@@
    The (big) mapping class group of a connected orientable surface which is either the 2-sphere or non-compact with genus either zero or infinite and whose end space is self-similar with a unique maximal end [14, Theorem 6.1].
@@
    A non-trivial closed subgroup of $GL_{n}(\mathbb{C})$ [22, Lemma 5.2].
@@
    A non-trivial connected locally compact group [22, Theorem 5.3].
@@
The subsequent lemma significantly simplifies the task of verifying the twisted Rokhlin property for a topological group.
@@
Note that the set $gV\phi(g)^{-1}$ is open whenever $V$ is open. Recall that, a topological group is Polish if its underlying topological space is separable and completely metrizable. The next result extends [14, Theorem 2.2] to the $\phi$-twisted case.
@@
It follows from [21, Corollary 5] that if $S$ is an infinite-type surface, then $\mathrm{MCG}(S)$ is a Polish group. Hence, $\mathrm{MCG}(S)$ has the $\phi$-twisted Rokhlin property if and only if it has the $\phi$-TJEP.
@@ 4. Twisted Rokhlin property of big mapping class groups
Recently, in [14, Theorem 6.1], Lanier and Vlamis have shown that the big mapping class group of a connected orientable infinite-type surface without boundary has the Rokhlin property if and only if the surface is non-compact whose genus is either zero or infinite and whose end space is self-similar with a unique maximal end.
@@ 4.1. When the surface has a unique maximal end
are called the positive and the negative parts of $\Sigma$, respectively. We now present an extension of [14, Proposition 4.2].
@@
Assertion (1) follows from [14, Proposition 4.2]. Let $c$ be a separating simple closed curve in $S$ such that $\mathfrak{t}(\Sigma_{c})=\Sigma_{c}$. Since $S$ is of infinite-type, we can choose another separating simple closed curve $c^{\prime}$ such that $\Sigma_{c}\cap\Sigma_{c^{\prime}}=\emptyset$, $\mathfrak{t}(\Sigma_{c^{\prime}})=\Sigma_{c^{\prime}}$ and $\Sigma_{c}\cong\Sigma_{c^{\prime}}$. Further, choose a separating simple closed curve $b$ such that the curves $\{c,c^{\prime},b\}$ bounds a pair of pants (a subsurface homeomorphic to the two-sphere with three disks removed) and $\{c,c^{\prime}\}\subset\Sigma_{b}$. Fix some hyperbolic metric on $\Sigma_{b}$ and choose an isometry $h:\Sigma_{b}\to\Sigma_{b}$ fixing $\partial(\Sigma_{b})$ identically such that $h(\Sigma_{c})=\Sigma_{c^{\prime}}$ and $h(\Sigma_{c}^{+})=\Sigma_{c^{\prime}}^{+}$. Define $f:S\to S$ by
@@
It follows from [14, Theorem 2.2 and Lemma 4.3] that $\mathrm{MCG}(S)$ has the Rokhlin property. Consequently, by Proposition 3.5, $\mathrm{MCG}(S)$ has the $\phi$-twisted Rokhlin property for each inner automorphism $\phi$ of $\mathrm{MCG}(S)$. By Proposition 2.4, we have $\operatorname{Out}(\mathrm{MCG}(S))\cong\mathbb{Z}_{2}$. Thus, in view of Proposition 3.5 and Proposition 3.10, it is enough to prove that $\mathrm{MCG}(S)$ has the $\hat{\mathfrak{t}}$-TJEP.
@@ 4.2. When the surface has exactly two maximal ends
It follows from [14, Lemma 5.4] that $\operatorname{Stab}(\mu)$ is a proper open normal subgroup of $\mathrm{MCG}(S)$. If $f\in h\operatorname{Stab}(\mu)\hat{\mathfrak{t}}(h)^{-1}$, then $f=hf^{\prime}\mathfrak{t}h^{-1}\mathfrak{t}$ for some $f^{\prime}\in\operatorname{Stab}(\mu)$. Since $\operatorname{Stab}(\mu)$ is normal, we see that
@@ 4.3. When the space of maximal ends is a Cantor space
This subsection deals with the case when every compact subsurface of $S$ is displaceable and $\mathcal{M}(S)$ is a Cantor space. The following result is a special case of [4, Theorem 14.1].
@@
By Proposition 3.5 and [14, Lemma 4.8], it is enough to prove that $\mathrm{MCG}(S)$ does not have the $\hat{\mathfrak{t}}$-twisted Rokhlin property. By results in [14, Proposition 4.6] (or [15, Lemma 4.18]), we choose a compact subsurface $\Sigma$ of $S$ such that
@@ 4.4. When the surface admits a compact non-displaceable subsurface
If $g=0$, then choose the simple closed curve $c$ to be as depicted in Figure 3 for $S_{0,6}$. Take $f=T_{c}$, where $c$ is the curve as above. By using Penner’s construction [4, Theorem 14.4], the homeomorphism $f\mathfrak{t}f\mathfrak{t}=T_{c}T_{\mathfrak{t}(c)}^{-1}$ is pseudo-Anosov since the collection $\{c,\mathfrak{t}(c)\}$ fills $S_{g,n}$. ∎
@@
In view of Proposition 3.5 and [14, Theorem 3.1], it is enough to prove that $\mathrm{MCG}(S)$ does not have the $\hat{\mathfrak{t}}$-twisted Rokhlin property. Let $\Sigma$ be a compact non-displaceable subsurface of $S$. By enlarging $\Sigma$, we can ensure that it is connected, has a stable Alexander system, satisfies $\mathfrak{t}(\Sigma)=\Sigma$ and its boundary $\partial(\Sigma)$ consists of at least six components with an even number of boundary components, each being essential and separating. Furthermore, $\Sigma$ remains a finite-type subsurface of $S$.

### 01 Jan 2024 v2

# Twisted Rokhlin property for mapping class groups

@@ Abstract.
In this paper, generalising the idea of the Rokhlin property, we explore the concept of the twisted Rokhlin property of topological groups. A topological group is said to exhibit the twisted Rokhlin property if, for each automorphism $\phi$ of the group, there exists a $\phi$-twisted conjugacy class that is dense in the group. We provide a complete classification of connected orientable infinite-type surfaces without boundaries whose mapping class groups possess the twisted Rokhlin property. Additionally, we prove that the mapping class groups of the remaining surfaces do not admit any dense $\phi$-twisted conjugacy class for any automorphism $\phi$. This supplements the recent work of Lanier and Vlamis on the Rokhlin property of big mapping class groups. We also prove that the mapping class group of each connected orientable infinite-type surface without boundary possesses the $R_{\infty}$-property.
@@
Big mapping class group; infinite type surface; $R_{\infty}$-property; Rokhlin property; twisted conjugacy; twisted Rokhlin property
@@ 1. Introduction
Given an automorphism $\phi$ of a group $G$, the map $(g,h)\mapsto gh\phi(g)^{-1}$, defines a left action of $G$ on itself. Two elements $x,y\in G$ are said to be $\phi$-twisted conjugate if they are in the same orbit under this action. The resulting orbits are referred as $\phi$-twisted conjugacy classes. The idea of twisted conjugacy arose from the work of Reidemeister [22]. It has deep connection with Nielsen fixed-point theory [6], and appears in Arthur–Selberg theory [24], algebraic geometry [12], and representation theory [10], to name a few. In recent years, there has been a tremendous amount of work on understanding twisted conjugacy in various classes of groups. See, for example [4, 8], for recent works on twisted conjugacy classes in mapping class groups and braid groups.
@@
Along the way, we also investigate the $R_{\infty}$-property of mapping class groups of surfaces and prove the following result.
Theorem 1.2.
Let $S$ be a connected orientable infinite-type surface without boundary. Then $\mathrm{MCG}(S)$ possesses the $R_{\infty}$-property.
@@ 2.1. Classification of surfaces
An end is called planar if it has a neighbourhood that can be embedded in the plane. Otherwise, it is called non-planar. Let $\mathcal{E}_{np}(S)$ be the subspace of $\mathcal{E}(S)$ consisting of non-planar ends. Clearly, $\mathcal{E}_{np}(S)$ is a closed set in $\mathcal{E}(S)$. The following theorem gives a classification of surfaces that we shall need [23, Theorem 1].
@@
Without loss of generality, we shall assume throughout that the surface $S$ is symmetric about the $xy$-plane when embedded in $\mathbb{R}^{3}$. Since $\mathcal{E}(S)$ is homeomorphic to a closed subset of the Cantor set, it can be embedded on the great circle of intersection of the $xy$-plane and the 2-sphere $\mathbb{S}^{2}$. Now, take a sequence of disjoint closed disks $D_{j}$ on $\mathbb{S}^{2}\setminus\mathcal{E}(S)$ which are symmetric about the $xy$-plane such that the set of accumulation points of this sequence is precisely $\mathcal{E}_{np}(S)$. We remove the interiors of the disks $D_{j}$ and suitably identify in pairs the boundary circles by orientation-reversing homeomorphisms such that the desired resulting surface is again symmetric about the $xy$-plane. See [23, Theorem 3] for the general construction. For instance, Figure 1 illustrates the construction of Loch Ness monster surface, where the infinite sequence of disjoint closed disks $\{D_{i}\}_{i\in\mathbb{N}}$ is chosen such that the radii of the disks decreases as it approaches a point of $\mathcal{E}(S)$.
@@
For a connected orientable surface $S$ (not necessarility of infinite-type), let $\operatorname{Homeo}^{+}(S)$ denote the group of orientation-preserving self-homeomorphisms of $S$. We define a binary relation $\preceq$ on the space of ends $\mathcal{E}(S)$ of $S$. For $x,y\in\mathcal{E}(S)$, we say that $y\preceq x$ if for every open neighbourhood $\mathcal{U}$ of $x$, there is an open neighbourhood $\mathcal{V}$ of $y$ and a homeomorphism $h\in\operatorname{Homeo}^{+}(S)$ such that $h(\mathcal{V})\subseteq\mathcal{U}$. An end $\mu$ is called maximal if, for any end $e$ satisfying $\mu\preceq e$, we must have $e\preceq\mu$. Let $\mathcal{M}(S)$ denote the set of all maximal ends of $S$. Note that for every end $e\in\mathcal{E}(S)$, there exists a maximal end $\mu\in\mathcal{M}(S)$ such that $e\preceq\mu$ [19, Proposition 4.7]. We say that two ends $e_{1}$ and $e_{2}$ are comparable if either $e_{1}\preceq e_{2}$ or $e_{2}\preceq e_{1}$. The space $\mathcal{E}(S)$ is said to be self-similar if given a decomposition $\mathcal{E}(S)=\mathcal{E}_{1}\sqcup\cdots\sqcup\mathcal{E}_{n}$ into a finite disjoint union of clopen subsets, there exists an index $i\in\{1,\ldots,n\}$ such that $\mathcal{E}_{i}$ contains an open homeomorphic copy of $\mathcal{E}(S)$. The idea of maximality and self-similarity for end spaces is due to [19].
@@ 2.2. Mapping class groups of surfaces
For a simple closed curve $c$ on a surface $S$, let $T_{c}$ denote the left-hand Dehn twist along $c$. By abuse of notation, we denote a curve and its isotopy class by the same symbol, and use the same convention for homeomorphisms as well. The following lemma, perhaps well-known, gives a relation between the Dehn twist and its conjugate by a homeomorphism of $S$. The proof of the following lemma is a slight modification of the arguments of [5, Fact 3.7]. See also [16, Lemma 2.1].
@@
is a basis for the topology on $\mathrm{MCG}(S)$. Note that, the center of the big mapping class group is always trivial [17, Proposition 2]. The following basic observation shall be used later.
@@ 2.3. Curve graphs of surfaces
We say that a simple closed curve $c$ on a surface $S$ is non-separating if $S\setminus c$ is connected; otherwise, it is called separating. The intersection number $i(c,c^{\prime})$ between two curves $c,c^{\prime}$ on $S$ is defined to be the minimum number of points of intersections between the representatives in their isotopy classes. A multicurve on $S$ is a set of pairwise (non-isotopic) disjoint curves.
The set $C(S)$ can be viewed as a graph whose vertex set is $C(S)$ and there is an edge between two vertices if the corresponding curves on $S$ are disjoint. The graph $C(S)$ is referred as curve graph of $S$. There is a natural metric $d_{C(S)}$ on $C(S)$ by defining the length of each edge as 1. We shall need the following result [20, Proposition 4.6].
@@
[21, Lemma 2.2] If $c\in\mathcal{C}(S)$ and $b,b^{\prime}\in\mathcal{C}(\Sigma)$ are two projections of $c$, then $d_{\mathcal{C}(\Sigma)}\left(b,b^{\prime}\right)\leq 2$.
@@ 3. Twisted Rokhlin property
    The homeomorphism group of the Cantor set [11, Theorem 2.6].
@@
    The group of orientation-preserving homeomorphisms of an even-dimensional sphere [11, Theorem 3.4].
@@
    The (big) mapping class group of a connected orientable surface which is either the 2-sphere or non-compact with genus either zero or infinite and whose end space is self-similar with a unique maximal end [18, Theorem 6.1].
@@
    A non-trivial closed subgroup of $GL_{n}(\mathbb{C})$ [26, Lemma 5.2].
@@
    A non-trivial connected locally compact group [26, Theorem 5.3].
@@
The subsequent proposition significantly simplifies the task of verifying the twisted Rokhlin property for a topological group.
@@
Note that the set $gV\phi(g)^{-1}$ is open whenever $V$ is open. Recall that, a topological group is Polish if its underlying topological space is separable and completely metrizable. The next result extends [18, Theorem 2.2] to the $\phi$-twisted case.
@@
It follows from [25, Corollary 5] that if $S$ is an infinite-type surface, then $\mathrm{MCG}(S)$ is a Polish group. Hence, $\mathrm{MCG}(S)$ has the $\phi$-twisted Rokhlin property if and only if it has the $\phi$-TJEP.
@@ 4. Twisted Rokhlin property of big mapping class groups
Recently, in [18, Theorem 6.1], Lanier and Vlamis have shown that the big mapping class group of a connected orientable infinite-type surface without boundary has the Rokhlin property if and only if the surface is non-compact whose genus is either zero or infinite and whose end space is self-similar with a unique maximal end.
@@ 4.1. When the surface has a unique maximal end
are called the positive and the negative parts of $\Sigma$, respectively. We now present an extension of [18, Proposition 4.2].
@@
Assertion (1) follows from [18, Proposition 4.2]. Let $c$ be a separating simple closed curve in $S$ such that $\mathfrak{t}(\Sigma_{c})=\Sigma_{c}$. Since $S$ is of infinite-type, we can choose another separating simple closed curve $c^{\prime}$ such that $\Sigma_{c}\cap\Sigma_{c^{\prime}}=\emptyset$, $\mathfrak{t}(\Sigma_{c^{\prime}})=\Sigma_{c^{\prime}}$ and $\Sigma_{c}\cong\Sigma_{c^{\prime}}$. Further, choose a separating simple closed curve $b$ such that the curves $\{c,c^{\prime},b\}$ bounds a pair of pants (a subsurface homeomorphic to the two-sphere with three disks removed) and $\{c,c^{\prime}\}\subset\Sigma_{b}$. Fix some hyperbolic metric on $\Sigma_{b}$ and choose an isometry $h:\Sigma_{b}\to\Sigma_{b}$ fixing $\partial(\Sigma_{b})$ identically such that $h(\Sigma_{c})=\Sigma_{c^{\prime}}$ and $h(\Sigma_{c}^{+})=\Sigma_{c^{\prime}}^{+}$. Define $f:S\to S$ by
@@
It follows from [18, Theorem 2.2 and Lemma 4.3] that $\mathrm{MCG}(S)$ has the Rokhlin property. Consequently, by Proposition 3.5, $\mathrm{MCG}(S)$ has the $\phi$-twisted Rokhlin property for each inner automorphism $\phi$ of $\mathrm{MCG}(S)$. By Proposition 2.4, we have $\operatorname{Out}(\mathrm{MCG}(S))\cong\mathbb{Z}_{2}$. Thus, in view of Proposition 3.5 and Proposition 3.10, it is enough to prove that $\mathrm{MCG}(S)$ has the $\hat{\mathfrak{t}}$-TJEP.
@@ 4.2. When the surface has exactly two maximal ends
It follows from [18, Lemma 5.4] that $\operatorname{Stab}(\mu)$ is a proper open normal subgroup of $\mathrm{MCG}(S)$. If $f\in h\operatorname{Stab}(\mu)\hat{\mathfrak{t}}(h)^{-1}$, then $f=hf^{\prime}\mathfrak{t}h^{-1}\mathfrak{t}$ for some $f^{\prime}\in\operatorname{Stab}(\mu)$. Since $\operatorname{Stab}(\mu)$ is normal, we see that
@@ 4.3. When the space of maximal ends is a Cantor space
This subsection deals with the case when every compact subsurface of $S$ is displaceable and $\mathcal{M}(S)$ is a Cantor space. The following result is a special case of [5, Theorem 14.1].
@@
By Proposition 3.5 and [18, Lemma 4.8], it is enough to prove that $\mathrm{MCG}(S)$ does not have the $\hat{\mathfrak{t}}$-twisted Rokhlin property. By results in [18, Proposition 4.6] (or [19, Lemma 4.18]), we choose a compact subsurface $\Sigma$ of $S$ such that
@@ 4.4. When the surface admits a compact non-displaceable subsurface
If $g=0$, then choose the simple closed curve $c$ to be as depicted in Figure 3 for $S_{0,6}$. Take $f=T_{c}$, where $c$ is the curve as above. By using Penner’s construction [5, Theorem 14.4], the homeomorphism $f\mathfrak{t}f\mathfrak{t}=T_{c}T_{\mathfrak{t}(c)}^{-1}$ is pseudo-Anosov since the collection $\{c,\mathfrak{t}(c)\}$ fills $S_{g,n}$. ∎
@@
In view of Proposition 3.5 and [18, Theorem 3.1], it is enough to prove that $\mathrm{MCG}(S)$ does not have the $\hat{\mathfrak{t}}$-twisted Rokhlin property. Let $\Sigma$ be a compact non-displaceable subsurface of $S$. By enlarging $\Sigma$, we can ensure that it is connected, has a stable Alexander system, satisfies $\mathfrak{t}(\Sigma)=\Sigma$ and its boundary $\partial(\Sigma)$ consists of at least six components with an even number of boundary components, each being essential and separating. Furthermore, $\Sigma$ remains a finite-type subsurface of $S$.